how much heat is absorbed during photosynthesis when 9.22 grams of glucose is produced
1
Expert's answer
2018-02-28T07:48:53-0500
To answer this question, one should know the formula of glucose as well as
photosynthesis net reaction:
6CO2 + 6H2O+ 2800 kJ  = C6H12O6 +6O2
Calculate number of moles of glucose
9.22g/180 g/mole = 0.051 moles;
The amount of heat will be calculated as
( 2800 kJ/mole) *0.051 mole = 143 kJ